A uniform meter rule AB is balanced on a knife edge placed 58cm from A when a mass of 20.0Kg is placed 25cm from B. Calculate the mass of the rule and the force exerted on the rule by the knife edge


1
Expert's answer
2021-11-21T17:27:04-0500

"AB = 1m"

"g= 9.8\frac{m}{s^2}"

"\text{let }K\text{-point where the knife is located}"

"AK = 58cm = 0.58m"

"BK = AB-AK = 0.42m"

"\text{let }M\text{-point where the mass is located }m_1=20kg"

"BM =25cm = 0.25m"

"KM = BK-BM= 0.18m"

"\text{let point } O \text{ midpoint of the ruler}"

"m - \text{this is the mass of the ruler}"

"AO = \frac{AB}{2}=0.5m"

"KO = AK-AO = 0.08m"

"KO*mg=KM*m_1g"

"m= \frac{KM*m_1}{KO}=\frac{0.18*20}{0.08}=45kg"

"N = mg+m_1g= 637N"


"\text{Answer: }"

"\text{the mass of the rule }45kg"

"\text{the force exerted on the rule by the knife edge }637N"